OpenStudy (anonymous):

A rock is launched from a cannon its height, h(x), can be represented by a quadratic function in terms of time, x, in seconds. After 1 second, the rock is 186 feet in the air after 2 seconds, it is 344 in the air. Find the height, in feet,of the rock after 8 seconds in the air.

OpenStudy (anonymous):

(You are basically solving for a,b, and c and forming a new quadratic equation.

OpenStudy (anonymous):

@mathmale @Nurali

OpenStudy (anonymous):

something like this -a*x^2+v*x you have to find a and v

OpenStudy (anonymous):

here is the data you do know -a*1^2+v*1=186 -a*2^2+v*2=344 SOlve that system somehow.

OpenStudy (anonymous):

I got the answer already it worked :D hehehe

OpenStudy (anonymous):

See, I have the formula, but it takes an unnecessary amount of time to use. I was actually looking for a shorter way to solve this problem,and this may be it..

OpenStudy (mathmale):

@timo86m "I got the answer already it worked :D hehehe" is definitely not an appropriate comment when it is YOUR responsibility to help the other person find the correct answer herself.

OpenStudy (anonymous):

@mathmale I haven't given her the answer

OpenStudy (anonymous):

I am just letting her know that my theory for solving it works. Now it is up to her to apply it.

OpenStudy (mathmale):

This is basically a physics problem, but is also seen in Math as well. If it's Math, you were probably given a model for the height of an object thrown straight up in the air as a function of t. Right? One such form of that wuold be \[h(t)=v _{0}t-\frac{ 1 }{ 2 }g*t^2\] where g is the acceleration due to gravity. You are given two pairs of data (t, h) that satisfy this equation, and you'll need to find the values of the initial velocity (v0) and of the acceleration due to gravity (g).

OpenStudy (mathmale):

Note that this IS a quadratic formula. I'm tempted to "cheat" a little and let g=-32 ft/(sec^2), substitute the coordinates from either point, and calculate v0 (the initial velocity). What do you think?

OpenStudy (anonymous):

it is a quadratic

OpenStudy (anonymous):

I was not given a model, but a quadratic equation: ax^2+bx+c. Thanks for your help, I like discovering new ways to solve this problem.

OpenStudy (anonymous):

Maybe we start from the beginning? :) I know all about cannons :3 Hey @alex-linne we know that h(x) is quadratic, so yeah, let's do that, the height is h, and time in seconds is x. \[\Large h(x) = \color{red}ax^2 +\color{blue}bx + \color{green}c\]

OpenStudy (anonymous):

@PeterPan it is easier to assume the cannon is at 0 initially os make c=0 just my assumption.

OpenStudy (anonymous):

Alex my way is the best. :D

OpenStudy (anonymous):

All in good time, Timo :P There doesn't seem to be a reason to rush. So @alex-linne do we proceed?

OpenStudy (anonymous):

Better. Okay, we can assume that at the beginning, the height of the rock (who the heck shoots rocks from cannons) is zero. After all, at time zero, like, just before it's fired, the rock is still on the ground... right? :)

OpenStudy (anonymous):

yeess.

OpenStudy (anonymous):

Okay. In Maths language, that amounts to \[\Large h(0) = 0\] OR \[\Large h(0) = \color{red}a(0)^2 + \color{blue}b(0) + \color{green}c = \color{green}c = 0\] Following? Doing this gives you a reason for c being zero... instead of just assuming it. So... do you follow? ^.^

OpenStudy (anonymous):

not true cuzz you can be firing from a cliff. SO it still an assumption but it works.

OpenStudy (anonymous):

Right. Nothing is happening at this point, so is that why there are zeros?

OpenStudy (anonymous):

Nothing happening at time zero (x = 0). So that narrows down your quadratic equation to \[\large h(x) = \color{red}ax^2 + \color{blue}bx\]

OpenStudy (anonymous):

simply put it is an assumption we made and it is correct. Only if they where to have told us we firing from a cliff of height h then we can plug it in.

OpenStudy (anonymous):

ahh.

OpenStudy (mathmale):

@TIMO86M: may i suggest that we focus on solving this problem in the shortest and most effective possible way, and only then offer alternative points of view AFTER that?

OpenStudy (anonymous):

@alex-linne The key to these word-problems is to know how to translate the words from English to Maths-Language :D So, translating "Height is zero, when time is zero" gives h(0) = 0 What about "Height is 186 after 1 second" ?

OpenStudy (anonymous):

I got 186=a+b+c..

OpenStudy (anonymous):

Sure. But we already know that c = 0 :P

OpenStudy (anonymous):

or at least, assumed it :)

OpenStudy (anonymous):

So... simplify?

OpenStudy (anonymous):

ah.. .-.

OpenStudy (anonymous):

Just replace c with zero. Now, and from now on :D 186 = a + b + 0 = a+b 186 = a + b We leave it at that, for now... What about "Height is 344 after 2 seconds" ?

OpenStudy (anonymous):

344=4a+2b? :o

OpenStudy (anonymous):

Nicely done XD

OpenStudy (anonymous):

Okay. a+b = 186 4a + 2b = 344 A system of linear equations. Can you solve it now? ^.^

OpenStudy (anonymous):

Not really ..haha. I got a strange answer .-.

OpenStudy (anonymous):

Really? How strange? What did you get?

OpenStudy (anonymous):

Wait, nevermind, I did something wrong >.>

OpenStudy (anonymous):

Keep going... the answer's pretty nice-looking (mind you, whatever you get for a and b, those won't be the final answers yet )

OpenStudy (anonymous):

I'm stuck at 158=3a+b.. >__>

OpenStudy (anonymous):

You're going to do that, are you? I suggest you multiply the first equation by 2 first, before doing anything of the sort :P

OpenStudy (anonymous):

2a + 2b = 372

OpenStudy (anonymous):

Ah, and then subtract that from 344=4a+2b? :o

OpenStudy (anonymous):

Yup ^.^

OpenStudy (anonymous):

a=14,so then what?

OpenStudy (anonymous):

I don't think so :P

OpenStudy (anonymous):

2a + 2b = 372 4a + 2b = 344 (4a + 2b) - (2a + 2b) = 344 - 372 Careful now :)

OpenStudy (anonymous):

ehhh ha ha..?

OpenStudy (anonymous):

How did you get a = 14?

OpenStudy (anonymous):

Show steps. Then we'll see where you made the mistake.